Calculate the mole fractions of ethanol, C_2 H_5 O
Answer #1

mole fraction = mol of i /total mol

then

mol of C2H5OH = mass/MW = 50/46 = 1.086

mol of H2O = mass/MW = 250/18 = 13.889

Total = 1.086+13.889 = 14.975

mol fraction of Ethanol = 1.086/14.975 = 0.07252

mol frac of water = 1-0.07252 = 0.92748